# Cascades in transport and optical conductivity of Twisted Bilayer Graphene

### v2

@@ Abstract
Among the plethora of correlated states in magic angle TBG, the cascades in the Density of States (DOS) [1, 2, 3, 4] and in the inverse compressibility [5, 6, 7], and the resistive states [8, 9, 6, 10, 11, 12], are the signatures of the electronic correlations that remain up to the highest temperatures. Experimentally, maxima in the resistivity as a function of doping appear around integer fillings up to temperatures much higher than those at which insulating behavior (resistivity increasing as the temperature decreases) is observed and, in many cases, even when the latter is absent  [8, 9, 6, 10, 11, 12]. Evidence of symmetry breaking order has been reported only up to a few kelvin. In many strongly correlated materials, Mott-like physics produces a redistribution of the spectral weight, even if the correlated material is in its normal state, above any symmetry-breaking ordering, or it is a metal [13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. However, due to the intrinsic topology of the TBG flat bands, most explanations of the correlated states observed in these systems, including those resilient with temperature, involved the assumption of symmetry breaking [2, 1, 5, 23, 24, 25, 26, 27].
The dichotomy between the local and topological properties can be incorporated within heavy fermion- like descriptions which include the remote bands [18, 28, 29, 30, 19]. In the different heavy fermion models for TBG, strongly correlated p+ and p- orbitals centered in the AA region of the moiré unit cell, AAp orbitals in the following, account for the most part of the spectral weight of the flat bands and are coupled to more itinerant orbitals only around $\Gamma$. Once the effect of the interactions are properly incorporated, the AAp orbitals tend to form local moments or heavy quasiparticles, depending on their filling [19, 20, 21, 22]. The electronic spectral weight $A({\bf k},\omega)$ obtained from DMFT calculations shows a characteristic momentum-selective incoherence and clear signatures of the Hubbard bands [19, 20] that could be studied in the future with nanoARPES or the Quantum Twisting Microscope [31, 32, 33]. Due to the intrinsic topology of the flat bands of TBG, the spectrum differs from that of standard Hubbard models and Mott insulators [19]. Without the need to involve any symmetry-breaking ordering, in the normal state, the spectral weight redistribution induced by the correlations produces a cascade spectrum in the DOS and asymmetric peaks in the inverse compressibility [19, 20, 21, 22] consistent with those observed experimentally  [1, 2, 3, 4, 5, 6, 7]. Little is known about the transport properties of TBG within this heavy fermion framework and the type of electrons governing them [34, 11].
@@
Figure 1: (a) Doping and energy dependent DOS, (b) inverse compressibility and (c) dc conductivity of TBG versus doping $\nu$. (d) Spectral weight at the Fermi level at CNP within the first Brillouin zone showing small pockets close to $\Gamma$.(e) Zoom of the low energy DOS for $\nu=1$ at temperatures T=5.8 K and T=0.7 K. (f) DOS at the Fermi level versus doping, including the total DOS (black), the AAp DOS (orange) and the itinerant orbitals DOS (blue). The quantities displayed in (a)-(f) have been obtained within the DMFT+H calculations. (g) Non-interacting band structure corresponding to the 1.08o TBG studied.
@@
We focus on a slightly particle-hole asymmetric TBG with twist angle $\theta=1.08^{\circ}$ and interlayer tunneling ratio between the AA and AB regions of the moiré unit cell $w_{0}/w_{1}=$0.7, Fig. 1(g) [45, 46, 47]. We fit the band structure with an 8 orbital model per valley and spin [48, 28, 19, 45] with topologically fragile flat bands. To calculate the interactions between the orbitals in the effective model, we assume a $1/r$ interaction between the electrons at the carbon atoms and relative dielectric constant $\epsilon=16$ [45]. We distinguish two types of orbitals: local AAp and itinerant $lc$. The two AAp orbitals, with $p_{+}$ and $p_{-}$ symmetries, centered at the AA region of the unit cell account for the spectral weight of the flat bands everywhere except close to the $\Gamma$ point. The intra-unit cell interaction between these local AAp orbitals, $U\sim$ 32.6 meV, is much larger than their bandwidth. These orbitals are strongly correlated [28, 19] and tend to form local moments. The $lc$ orbitals are less correlated. The model resembles an extended heavy-fermion like model with a strongly ${\bf k}$-dependent hybridization between the local AAp and the itinerant $lc$ orbitals, concentrated around $\Gamma$, see Fig. S1 [45] and [18, 28, 29, 19, 49, 50, 21, 20, 51]. The interactions involving the itinerant orbitals are of order $U/2$ and cannot be neglected [45, 28, 19, 1, 29]. As in our previous work, we treat the intra and inter-orbital onsite interaction $U$ between the AAp orbitals within DMFT and the other interactions at a mean field Hartree level, within a double self-consistency loop (DMFT+H), see [19, 45]. Temperature is $5.8\,\rm K$ except otherwise indicated. Spontaneous symmetry-breaking is not allowed in the calculation.
@@
A strong resetting behavior with minima in the dc conductivity is also seen in Fig. 1(c). These minima around integers produce resistive peaks as those found in transport measurements up to tens of kelvin  [8, 9, 6, 10, 11, 12]. Our calculations hence show that the presence of these peaks in the resistivity does not require symmetry-breaking, forbidden in the computation. At the Charge Neutrality Point (CNP), the conductivity is notably enhanced with respect to the one at the other integers and to the one expected for a Dirac semimetal. This is a consequence of the presence of compensated electron and hole pockets close to $\Gamma$, Fig. 1(d). These pockets emerge due to the hybridization between the itinerant electrons at $\Gamma$ and the local AAp, whose spectral weight is shifted to the Hubbard bands at higher energies, Fig. 3(d). The suppression of the DOS at non-zero integers was previously interpreted by us[19] and others [3, 20] as a pseudogap. A recent work has studied the formation of decoupled moments in isolated topological bands and related it to the TBG flat bands [52] and finds hard gaps at non-zero integers. In Fig. 1(f) we compare the DOS close to the Fermi level at $\nu=1$ at T=5.8 K, the temperature of the calculations, and at T=0.7 K, a temperature at which the incoherent spectral weight is strongly suppressed. The numerical data at these temperatures seem more consistent with an asymmetric nodal gap or a pseudogap than with a hard gap, but we cannot rule out the latter appearing at larger interactions.
@@
Figure 3: (a) DMFT+H optical conductivity showing resets in the intensity at integer fillings, oscillations in the frequency at which the intensity peaks. (b)-(c) Optical spectrum respectively in the absence of interactions and in the Hartree approximation.(d) Spectral weight $A({\bf{k}},w)$ from the DMFT+H calculations at doping $\nu=0,0.21,0.5,1.05,1.88~{}\rm{and}~{}3.5$ as a function of momentum; and (e) Spectral weight $A({\bf{k}},w)$ as a function of doping $\nu$ at fixed momenta: From left to right $K$, along $K-\Gamma$ direction, $\Gamma$, and $M$, marked in Fig. 1(d). The colorbar in (d) is also valid for (e).
@@
Fig. S1(a) shows the non-interacting band structure for one of the valleys, obtained from the continuum model (black) for a 1.08 TBG with vF =0.87$\times$106 m/s, interlayer tunneling ratio between the AA and AB regions $w_{0}/w_{1}=0.7$ and AB tunneling amplitude $w_{1}$= 110 meV, and the effective 8-orbital fitting used (blue) [46, 47, 48]. In the continuum model we have kept the $\sin(\theta/2)$ term. This results in slightly particle hole asymmetric bands. The flat bands width at the $\Gamma$ point is 7.7 meV and 1.4 meV at M. The onset of the remote conduction and valence bands is respectively at 35.5 and 37 meV.
@@
The interactions between the moiré orbitals are calculated projecting the moiré Wannier functions onto the two layers atomic sublattices and assuming a $1/r$ decay of the interaction between the electrons in the carbon atoms and a relative dielectric constant $\epsilon=16$. We do not find significant differences between the density-density interactions in the different valleys. The intra and inter-orbital density-density interactions used in the self-consistent calculation are shown in Fig. S1(b). The onsite AAs-AAp interaction has been reduced a 26 % from the computed value to decrease the small particle-hole asymmetry induced by the 8- orbital fitting when the interactions are included. The inter-site density-density interactions are kept up to a distance of 7 neighboring moiré unit cells in the self-consistent calculation and to compute the effect of the gate in the chemical potential. Exchange, Hund and pair hopping terms are much smaller than the density-density terms and we neglect them here [28].
To avoid double-counting the effect of the interactions we subtract the Hartree contribution at the CNP [54] produced by all the interactions, including $U$. For finite doping the energy cost to charge the unit cell is compensated by the gate, whose contribution $V_{g}$ is subtracted from the chemical potential to calculate the inverse compressibility. We approximate $V_{g}$ as the average energy required to add $\nu$ electrons homogeneously to the 8 orbitals included in the model.
@@
The impact of the interactions on the spectral weight is calculated with a double self-consistent DMFT+Hartree loop, as detailed in [19], using the CTQMC impurity solver developed in [59, 60]. The analytic continuation is performed with a maximum entropy method [61] as implemented in [60]. The value of $\Gamma_{\rm{AA}_{p}}$ in Fig. 2 is computed at $\omega=0$ after doing the analytic continuation.
[…]
@@
×$\times\mathbb{A}(\omega,{\bf k})\mathbb{V}_{\alpha}({\bf k})\bigg{]}\frac{f(\omega)-f(\omega+\Omega)}{\Omega}.$
@@
Figs. S4 and S5 show the evolution of the DMFT+H bandstructure for many different dopings in the range $-4<\nu<4$, which allows us to see the alternance of highly incoherent bands above the integers, evolving towards more coherent quasiparticle bands at the Fermi level below the integers. The incoherent spectral weight is spread over a wide range of energies. The formation of prominent Hubbard bands above the integers indicates the formation of local moments. With doping their width and intensity evolve, as spectral weight is transferred to form the quasiparticle. The Hubbard bands are bended towards $\Gamma$ due to the hybridization between the local and the itinerant electrons. At nonzero integer fillings the DOS at the Fermi level is suppressed, but the still highly coherent, flat band remains very close to the Fermi level. Above the integer this band shifts away from the Fermi level becoming very incoherent and giving rise to a new Hubbard band.
@@
We have excluded from the fittings in Fig. 2 the dopings very close to the CNP (-0.14 $<\nu<$ 0.14) and the integer dopings. From visual inspection, a Drude fitting is not valid for these dopings. In the case of the integers this is due to the vanishing conductivity at zero frequency. Close to the CNP the shape is not Drude-like due to the contribution of very low frequency interband transitions and the strong incoherence.

### v3

@@ Abstract
Among the plethora of correlated states in magic angle TBG, the cascades in the Density of States (DOS) [59, 12, 32, 60] and in the inverse compressibility [63, 51, 50], and the resistive states [7, 46, 51, 47, 41, 16], are the signatures of the electronic correlations that remain up to the highest temperatures. Experimentally, maxima in the resistivity as a function of doping appear around integer fillings. Insulating behavior, i.e. resistivity increasing as temperature decreases, and evidence of symmetry breaking order, have been detected only up to few kelvin. On the other hand, the maxima in the resistivity around integer fillings are present up to much higher temperatures and even when the low temperature insulating behavior is absent  [7, 46, 51, 47, 41, 16]. In many strongly correlated materials, Mott-like physics produces a redistribution of the spectral weight. Mott effects remain at temperatures above symmetry-breaking orderings, when the correlated material is in its normal state, and do not necessarily require insulating behavior [28, 2, 15, 20, 45, 22, 14, 49, 26, 34, 24]. Due to the intrinsic topology of the TBG flat bands, most explanations of the correlated states observed in TBG, including those resilient with temperature, have typically involved the assumption of symmetry breaking [12, 59, 63, 17, 9, 53, 36, 11].
The dichotomy between the local and the topological properties can be incorporated within heavy fermion- like descriptions which include the remote bands [22, 4, 55, 54, 14]. These models deal with extended moiré orbitals, not atomic orbitals. In the different heavy fermion models for TBG, strongly correlated p+ and p- orbitals located in the center of the moiré unit cell, the AA region, account for the most part of the spectral weight of the flat bands. These orbitals, named AAp orbitals in the following, are coupled to more itinerant orbitals only around $\Gamma$. Once the effect of the interactions are properly incorporated, the AAp orbitals tend to form local moments or heavy quasiparticles, depending on their filling [14, 49, 26, 34]. The electronic spectral weight $A({\bf k},\omega)$ obtained from DMFT calculations shows a characteristic momentum-selective incoherence and clear signatures of the Hubbard bands [14, 49] that could be studied in the future with nanoARPES or the Quantum Twisting Microscope [40, 39, 29]. Due to the intrinsic topology of the flat bands of TBG, the spectrum differs from that of standard Hubbard models and Mott insulators [14]. Without the need to involve any symmetry-breaking ordering, in the normal state, the spectral weight redistribution induced by the correlations produces a cascade spectrum in the DOS and asymmetric peaks in the inverse compressibility [14, 49, 26, 34] consistent with those observed experimentally  [59, 12, 32, 60, 63, 51, 50]. Little is known about the transport properties of TBG within this heavy fermion framework and the type of electrons governing them [6, 41].
@@
Figure 1: (a) Doping and energy dependent DOS, (b) inverse compressibility and (c) dc conductivity of TBG versus doping $\nu$. (d) Spectral weight at the Fermi level at CNP within the first Brillouin zone showing small pockets close to $\Gamma$. We highlight momenta $\Gamma$, M, K and kΓ-K at which the doping dependence spectral weight is plotted in Fig. 3(e). (e) Zoom of the low energy DOS for $\nu=1$ at temperatures T=5.8 K and T=0.7 K. (f) DOS at the Fermi level versus doping, including the total DOS (black), the AAp DOS (orange) and the itinerant orbitals DOS (blue). The quantities displayed in (a)-(f) have been obtained within the DMFT+H calculations. (g) Non-interacting band structure corresponding to the 1.08o TBG studied.
@@
We focus on a slightly particle-hole asymmetric TBG with twist angle $\theta=1.08^{\circ}$ and interlayer tunneling ratio between the AA (center) and the AB (honeycomb vertex) regions of the moiré unit cell $w_{0}/w_{1}=$0.7, Fig. 1(g) [30, 3, 44]. We fit the band structure with an 8 orbital model per valley and spin, see description in the Supplementary Information [8, 4, 14, 30], with topologically fragile flat bands. We distinguish two types of orbitals: two local AAp and six itinerant $lc$ orbitals [30]. The two AAp orbitals, with $p_{+}$ and $p_{-}$ symmetries, centered at the AA region of the unit cell, account for the spectral weight of the flat bands everywhere except close to the $\Gamma$ point. To calculate the interactions between the orbitals in the effective model, we assume a $1/r$ interaction between the electrons at the carbon atoms and relative dielectric constant $\epsilon=16$ [30]. The intra-unit cell interaction between the local AAp orbitals, $U\sim$ 32.6 meV, is much larger than their bandwidth. These orbitals are strongly correlated [4, 14] and tend to form local moments. The $lc$ orbitals are less correlated. The model resembles an extended heavy-fermion like model with a strongly $\bf k$}-dependent hybridization between the local AAp and the itinerant $lc$ orbitals, concentrated around $\Gamma$ [22, 4, 55, 14, 13, 25, 26, 49, 62], see Fig. S1 and the Supplementary Information [30]. Different to more standard heavy fermion models, the interactions involving the itinerant orbitals are of order $U/2$ and cannot be neglected [30, 4, 14, 59, 55]. As in our previous work, we treat the intra and inter-orbital onsite interaction $U$ between the AAp orbitals within DMFT and the other interactions at a mean field Hartree level, within a double self-consistency loop (DMFT+H), see [14, 30]. Temperature is $5.8\,\rm K$ except otherwise indicated. Spontaneous symmetry-breaking is not allowed in the calculation.
@@
A strong resetting behavior with minima in the dc conductivity is also seen in Fig. 1(c). These minima around integers produce resistive peaks as those found in transport measurements up to tens of kelvin  [7, 46, 51, 47, 41, 16]. Our calculations hence show that the presence of these peaks in the resistivity does not require symmetry-breaking, forbidden in the computation. At the Charge Neutrality Point (CNP), the conductivity is notably enhanced with respect to the one at the other integers and to the one expected for a Dirac semimetal. This is a consequence of the presence of compensated electron and hole pockets close to $\Gamma$, Fig. 1(d). These pockets emerge due to the hybridization between the itinerant electrons at $\Gamma$ and the local AAp, whose spectral weight is shifted to the Hubbard bands at higher energies, Fig. 3(d). The suppression of the DOS at non-zero integers was previously interpreted by us[14] and others [32, 49] as a pseudogap. A recent work has studied the formation of decoupled moments in the isolated TBG flat bands [37]. At non-zero integers, they find hard gaps, i.e. the full Fermi surface is gapped. In Fig. 1(f) we compare the DOS close to the Fermi level at $\nu=1$ at T=5.8 K, the temperature of the calculations, and at T=0.7 K, a temperature at which the incoherent spectral weight is strongly suppressed. The numerical data at these temperatures seem more consistent with an asymmetric nodal gap or a pseudogap than with a hard gap, but we cannot rule out the latter appearing at larger interactions.
@@
Figure 3: (a) DMFT+H optical conductivity showing resets in the intensity at integer fillings, oscillations in the frequency at which the intensity peaks. (b)-(c) Optical spectrum respectively in the absence of interactions and in the Hartree approximation.(d) Spectral weight $A({\bf{k}},w)$ from the DMFT+H calculations at doping $\nu=0,0.21,0.5,1.05,1.88~\rm{and}~3.5$ as a function of momentum; and (e) Spectral weight $A({\bf{k}},w)$ as a function of doping $\nu$ at fixed momenta. From left to right: $K$, the kk-Γ point situated along the $K-\Gamma$ direction, $\Gamma$, and $M$, all of them marked in Fig. 1(d). The colorbar in (d) is also valid for (e).
@@
[…]
@@
The interactions between the moiré orbitals are calculated projecting the moiré Wannier functions onto the two layers atomic sublattices and assuming a $1/r$ decay of the interaction between the electrons in the carbon atoms and a relative dielectric constant $\epsilon=16$. We do not find significant differences between the density-density interactions in the different valleys. The intra and inter-orbital density-density interactions used in the self-consistent calculation are shown in Fig. S1(b). The onsite AAs-AAp interaction has been reduced a 26 % from the computed value to decrease the small particle-hole asymmetry induced by the 8- orbital fitting when the interactions are included. The inter-site density-density interactions are kept up to a distance of 7 neighboring moiré unit cells in the self-consistent calculation and to compute the effect of the gate in the chemical potential. Exchange, Hund and pair hopping terms are much smaller than the density-density terms and we neglect them here [4]. Specifically the onsite Hund’s and pair hopping terms are of order 10-3U. For instance, the intravalley Hund’s coupling is suppressed because the product of the envelope of the Wannier function changes sign within the unit cell while for the intervalley case the suppresion originates in product of Bloch factors oscillating at the atomic scale which integrate to a negligible value [4]. The small overlap between the Wannier functions in different unit cells suppresses the exchange and assisted hopping terms, of order 10-2U, between local orbitals [4]. The exchange between the local and itinerant orbitals may help selecting a specific symmetry breaking order, but it does affect the normal state physics discussed here. To avoid double-counting the effect of the interactions we subtract the Hartree contribution at the CNP [19] produced by all the interactions, including $U$. For finite doping the energy cost to charge the unit cell is compensated by the gate, whose contribution $V_{g}$ is subtracted from the chemical potential to calculate the inverse compressibility. We approximate $V_{g}$ as the average energy required to add $\nu$ electrons homogeneously to the 8 orbitals included in the model.
@@
The impact of the interactions on the spectral weight is calculated with a double self-consistent DMFT+Hartree loop [14]. We use single-site DMFT to deal with the onsite interaction between the local orbitals and treat all the other interactions at the Hartree level. To solve the DMFT calculations we use the continuous time quantum Monte Carlo (CTQMC) impurity solver implemented in  [21, 33]. In the Hartree approximation the interactions produce a modification of the onsite potential of each orbital. These onsite orbitals depend on the orbital fillings. The outcome of the DMFT calculation is the self-energy of the local orbitals. We prohibit symmetry breaking imposing that this self-energy is diagonal, equal for the two p+ and p- orbitals and valley and spin independent. The coupling to the other orbitals as well as the effect of the Hartree potentials enter in the DMFT calculation through a $\Delta_{0}$ hybridization function. A number of DMFT iterations at fixed DMFT chemical potential are run starting from a given set of Hartree onsite potentials. We run a number of DMFT iterations for the local orbitals for a fixed DMFT chemical potential. Using the DMFT self-energy we recalculate the fillings of the local and itinerant orbitals. With the new fillings we calculate new Hartree onsite potentials and then run new DMFT iterations without changing the DMFT chemical potential. We continue doing Hartree and DMFT loops until convergence is reached for both the orbital fillings and self-energy. The calculation is done at finite temperature and the self-energy obtained is in Matsubara frequencies. The analytic continuation is performed with a maximum entropy method [31] as implemented in [33]. The value of $\Gamma_{\rm{AA}_{p}}$ in Fig. 2 is computed at $\omega=0$ after doing the analytic continuation.
Figure S2: (a) and (b) Band structure for the 1.08 and $w_{0}/w_{1}=0.7$ twist angle TBG in the Hartree approximation corresponding to $\nu$=1.0 and $\nu$=3.5 calculated using the same interactions as in the DMFT+H calculations. (c), (e) and (g) Orbital resolved density of states for AAp and $\it lc$ orbitals, dc conductivity and low frequency optical conductivity, respectively, as a function of doping $\nu$ in the absence of interactions. (d), (f) and (h) Same in the Hartree approximation.
@@
$\times\mathbb{A}(\omega,{\bf k})\mathbb{V}_{\alpha}({\bf k})\bigg]\frac{f(\omega)-f(\omega+\Omega)}{\Omega}.$
@@
Figs. S4 and S5 show the evolution of the DMFT+H bandstructure for many different dopings in the range $4<\nu<4$, which allows us to see the alternance of highly incoherent bands above the integers, evolving towards more coherent quasiparticle bands at the Fermi level below the integers. The incoherent spectral weight is spread over a wide range of energies. The formation of prominent Hubbard bands above the integers indicates the formation of local moments. With doping their width and intensity evolve, as spectral weight is transferred to form the quasiparticle. The Hubbard bands are bended towards $\Gamma$ due to the hybridization between the local and the itinerant electrons. At nonzero integer fillings the DOS at the Fermi level is suppressed, but the still highly coherent, flat band remains very close to the Fermi level. Above the integer this band shifts away from the Fermi level becoming very incoherent and giving rise to a new Hubbard band.
@@
We have excluded from the fittings in Fig. 2 the dopings very close to the CNP (-0.14 $\nu<$ 0.14) and the integer dopings. From visual inspection, a Drude fitting is not valid for these dopings. In the case of the integers this is due to the vanishing conductivity at zero frequency. Close to the CNP the shape is not Drude-like due to the contribution of very low frequency interband transitions and the strong incoherence.